Construct two composite functions

(g º f) (x)=
(g º f) (2)=
(f º g) (x)=
(f º g) (2)=

Answer by stanbon(75887)

Construct two composite functions:
f(x) = x + 5 ; g(x) = x^2
-------
(g º f) (x)= g[f(x)] = g[x+5] = (x+5) = x^2 + 10x + 25
-----------------------------------------------------
(g º f) (2)= 2^2 + 10*2 + 25 = 4+100+25 = 129
----------------------------------------------------
(f º g) (x)= f[g(x)] = f[x^2] = x^2+5
-------------------------------------------
(f º g) (2)= 2^2+5 = 4+5 = 9
==================================
Cheers,
Stan H.
=================================